OpenStudy (anonymous):

What is the inverse of y=-4x^2-2 and graph? I am not sure if I am doing this right.

OpenStudy (anonymous):

find x in terms of y 4x^2 = -2-y x^2 = (-2-y) / 4 x = sqrt [ (-2-y)] / 2 f-1(y) = sqrt[ -2-x]) / 2 the graph of the inverse of y is the reflection of y in the line y=x

OpenStudy (anonymous):

Ok thank you